# MuonEq: Balancing Before Orthogonalization with Lightweight Equilibration

### 30 Mar 2026 v1

@@ Appendix D Normalization and whitening
Let $\mathbf{S}_{t}:=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\gamma_{\mathrm{ns}}:=1-\varepsilon_{\mathrm{ns}},c_{m,n}^{\mathrm{ns}}:=\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}$. By Assumption 3.2 and the descent lemma, we have
$\leq f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2},$ (7)
where the last step uses Lemma G.3.
Again by Lemma G.3,
$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle$ $=\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\langle\mathbf{S}_{t},\mathbf{O}_{t}\rangle$
$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-\|\mathbf{S}_{t}\|_{F}\|\mathbf{O}_{t}\|_{F}$
$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\|\mathbf{S}_{t}\|_{F}$
$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\nabla f(\mathbf{X}_{t})\|_{F}-\Bigl(\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)\|\mathbf{S}_{t}\|_{F}.$ (8)
Substituting (H) into (H),
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+ac_{m,n}^{\mathrm{ns}}\eta_{t}\|\mathbf{S}_{t}\|_{F}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2}.$ (9)
@@
Hence
$f(\mathbf{X}_{t+1})\leq$ $\ f(\mathbf{X}_{t})-\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}$
$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2}.$
@@
$\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $\leq f(\mathbf{X}_{1})-f^{\star}+\frac{a}{2L}\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\sum_{t=1}^{T}\eta_{t}^{2}.$ (10)
A direct repetition of the proof of Lemma E.3, using $\|\mathbf{O}_{t}\|_{F}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n$ from Lemma G.3, gives
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{(1-h_{t+1})^{2}}{h_{t+1}}L^{2}\eta_{t}^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+h_{t+1}^{2}\sigma^{2}.$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr).$
Applying Lemma E.2 with
$E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},\qquad A_{t+1}=h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr),$
@@
$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)\sum_{t=1}^{T}\frac{1}{t+1}$
$\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)(1+\ln T).$
Hence, since $h_{t+1}\leq h_{t}$,
$\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)(1+\ln T).$
@@
Substituting these bounds into (H),
$\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}.$
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}}{\bigl(a\gamma_{\mathrm{ns}}/\sqrt{m}\bigr)T^{1/4}}.$
This completes the proof. ∎
Appendix I More Results
@@
(a) Step 1%
(b) Step 10%
(c) Step 50%
(d) Step 100%
@@
Appendix J Experimental Details
J.1 Pretraining on C4
@@
Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq
Max Learning Rate 5e-3 3e-3 3e-3 3e-3 3e-3
@@
Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq
Max Learning Rate 1.5e-3 2e-3 1.5e-3 2e-3 2e-3
@@
Gamma 0.025 0.05
[…]
J.2 Ablation Study Details
@@
Hyper-parameter SGD AdamW Muon-Nes MuonEq (RC) MuonEq (C) MuonEq (R)
Max Learning Rate 5e-2 1e-3 5e-2 5e-2 5e-2 5e-2
@@
Hyper-parameter AdamW Muon-Nes MuonEq (RC) MuonEq (C) MuonEq (R)
Max Learning Rate 2e-3 2e-3 2e-3 2e-3 2e-3
@@
We compare the three MuonEq variants RC, R, and C. RC applies two-sided row/column normalization throughout training, R applies row normalization throughout training, and C applies column normalization throughout training. We take R as the default variant. In all cases, the orthogonalization routine, optimizer states, and scalar hyperparameters are unchanged.
@@
This ablation is directly connected to Eq. (5). RC provides the strongest two-sided spectral correction and mainly targets the finite-step Newton–Schulz term. R and C test the two one-sided geometries under the same optimizer configuration. Since MuonEq targets hidden matrix weights, R is the more relevant default in our setting, while C is included as the column-sided companion form.
J.3 A Note on Row/Column Terminology in Prior Work
A careful comparison to prior row/column normalization papers requires fixing the matrix convention. [13] write linear weights as $\theta\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ and define column-wise normalization along the output dimension. Under the standard deep-learning storage layout $\mathbf{W}=\theta^{\top}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$, this corresponds to row normalization of the stored hidden-weight matrix. Their empirical observation that row-wise normalization can be unstable is also traced largely to the LM-head, where row-wise normalization produces extreme gradient values. This is not the same setting as MuonEq, which is aimed at hidden matrix weights.

### 10 May 2026 v2

@@ Appendix E Normalization and whitening
Let $\lambda_{0}:=\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}}.$ Since $\rho<a\gamma_{\mathrm{ns}}/\sqrt{m}\leq a/\sqrt{m}$, we have
$\lambda_{0}\leq\frac{1}{4(1+\varepsilon_{\mathrm{ns}})\sqrt{mn}}\leq 1.$
Hence
$0\leq\eta_{t}\lambda_{t}\leq\lambda_{0}t^{-1}\leq 1.$
Using Lemma H.4,
$\|\mathbf{O}_{t}\|_{F}\leq(1+\varepsilon_{\mathrm{ns}})\sqrt{n}.$
Therefore,
$\|\mathbf{X}_{t+1}\|_{F}$ $=\|(1-\eta_{t}\lambda_{t})\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t}\|_{F}$
$\leq(1-\eta_{t}\lambda_{t})\|\mathbf{X}_{t}\|_{F}+a\eta_{t}\|\mathbf{O}_{t}\|_{F}$
$\leq\|\mathbf{X}_{t}\|_{F}+a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\eta_{t}.$
By induction and $\sum_{s=1}^{t-1}s^{-3/4}\leq 4t^{1/4}$,
$\|\mathbf{X}_{t}\|_{F}\leq\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}t^{1/4}.$
Thus
$\lambda_{t}\|\mathbf{X}_{t}\|_{F}\leq\lambda_{0}t^{-1/4}\Bigl(\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}t^{1/4}\Bigr)\leq\rho.$
Finally,
$\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}$ $=\eta_{t}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}$
$\leq\eta_{t}\Bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\Bigr)$
This completes the proof. ∎
Appendix I Proofs of Theorem 3.5 and Corollary 3.6
Proof.
Let $\mathbf{S}_{t}:=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\gamma:=1-\varepsilon_{\mathrm{ns}},$ and define $c_{m,n}^{\mathrm{ns}}:=\frac{\gamma}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}.$
By (7), we have $\frac{a\gamma}{\sqrt{m}}-\rho>0$. By Assumption 3.2 and the descent lemma,
$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})+\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$ (8)
$=f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle-\lambda_{t}\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\rho\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{L}{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}\eta_{t}^{2},$
where the last step uses Lemma H.5 and Cauchy’s inequality.
Again by Lemma H.4, we have
$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle$ $=\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\langle\mathbf{S}_{t},\mathbf{O}_{t}\rangle$ (9)
$\geq\frac{\gamma}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-\|\mathbf{S}_{t}\|_{F}\|\mathbf{O}_{t}\|_{F}$
$\geq\frac{\gamma}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\|\mathbf{S}_{t}\|_{F}$
$\geq\frac{\gamma}{\sqrt{m}}\|\nabla f(\mathbf{X}_{t})\|_{F}-\Bigl(\frac{\gamma}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)\|\mathbf{S}_{t}\|_{F}.$
Substituting (9) into (8), we have
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\left(\frac{a\gamma}{\sqrt{m}}-\rho\right)\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+ac_{m,n}^{\mathrm{ns}}\eta_{t}\|\mathbf{S}_{t}\|_{F}+\frac{L}{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}\eta_{t}^{2}.$ (10)
@@
Let $d_{\rho}^{\mathrm{ns}}:=a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho$ and $\kappa_{\rho}^{\mathrm{ns}}:=\frac{a\gamma}{\sqrt{m}}-\rho.$ Hence
$f(\mathbf{X}_{t+1})\leq$ $f(\mathbf{X}_{t})-\kappa_{\rho}^{\mathrm{ns}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}$
$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{L}{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}\eta_{t}^{2}.$
@@
$\kappa_{\rho}^{\mathrm{ns}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $\leq f(\mathbf{X}_{1})-f^{\star}+\frac{a}{2L}\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ (11)
$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{L}{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}\sum_{t=1}^{T}\eta_{t}^{2}.$
By Lemma F.3, using $\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}\leq d_{\rho}^{\mathrm{ns}}\eta_{t}$ from Lemma H.5, we have
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{(1-h_{t+1})^{2}}{h_{t+1}}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}\eta_{t}^{2}+h_{t+1}^{2}\sigma^{2}.$
@@
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr).$
Applying Lemma F.2 with
$E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},\qquad A_{t+1}=h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr),$
@@
$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)\sum_{t=1}^{T}\frac{1}{t+1}$
$\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)(1+\ln T).$
Hence, since $h_{t+1}\leq h_{t}$, we have
$\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)(1+\ln T).$
@@
Substituting the above bounds into (11),
$\kappa_{\rho}^{\mathrm{ns}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ns}}(1+\ln T)+C_{2,\rho}^{\mathrm{ns}}.$
where
$C_{1,\rho}^{\mathrm{ns}}:=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)+aL\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2},C_{2,\rho}^{\mathrm{ns}}:=\frac{a\sigma^{2}}{L}+\frac{3L}{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}.$
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ns}}(1+\ln T)+C_{2,\rho}^{\mathrm{ns}}}{\kappa_{\rho}^{\mathrm{ns}}\cdot T^{1/4}}.$
Setting $\varepsilon_{\mathrm{ns}}$ to zero yields the corresponding exact-polar result for Theorem 3.5 with decoupled weight decay. Setting additionally $\rho=0$ recovers the zero weight decay proof. This completes the proof. ∎
Appendix J More Results
[…]
((a))
((b))
@@
Appendix K Experimental Details
K.1 Pretraining on C4
[…]
$\mathbf{M}_{t}^{\rm Nes}=\beta\mathbf{M}_{t-1}^{\rm Nes}+(1-\beta)\mathbf{G}_{t}+\gamma\beta\bigl(\mathbf{G}_{t}-\mathbf{G}_{t-1}\bigr),\qquad\mathbf{G}_{t-1}:=\nabla f(\mathbf{X}_{t-1};\xi_{t-1}).$
The MuonEq-Nes convention used in Algorithm 1 corresponds to $\gamma=1-\beta$. Indeed, if the EMA buffer is $\mathbf{M}_{t}^{\rm EMA}=\beta\mathbf{M}_{t-1}^{\rm EMA}+(1-\beta)\mathbf{G}_{t}$ and the Nesterov input is $\widetilde{\mathbf{M}}_{t}=\beta\mathbf{M}_{t}^{\rm EMA}+(1-\beta)\mathbf{G}_{t}$, then $\widetilde{\mathbf{M}}_{t}$ satisfies the recurrence above with $\gamma=1-\beta$. Thus the reported Muon momentum $\beta=0.95$ corresponds to $\gamma=0.05$; the actual coefficient multiplying $\mathbf{G}_{t}-\mathbf{G}_{t-1}$ is $\gamma\beta=0.0475$.
((a))
((b))
((c))
((d))
@@
((a))
((b))
((c))
((d))
Figure 8: Validation perplexity heatmaps on GPT2-small/FineWeb (2.6B tokens) over learning rate and momentum at $K=4$ and $K=5$.
@@
Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq MuonEq-Nes
Max Learning Rate 5e-3 3e-3 3e-3 3e-3 3e-3 3e-3
@@
Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq MuonEq-Nes
Max Learning Rate 1.5e-3 2e-3 1.5e-3 2e-3 2e-3 2e-3
@@
Gamma 0.025 0.05 0.05
Table 6: Hyperparameters used for training LLaMA2-1B on C4
Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq MuonEq-Nes
Max Learning Rate {5e-4, 8e-4}
Warmup Steps 1000
Batch Size 512
Maximum Length 4096
Weight Decay 0.1
$(\beta_{1},\beta_{2})$ (0.9,0.95)
Muon-Momentum 0.95
Gamma 0.025 0.05 0.05
K.2 Ablation Study Details
@@
Hyper-parameter SGD AdamW Muon Muon-Nes MuonEq-Nes (RC) MuonEq-Nes (C) MuonEq-Nes (R)
Max Learning Rate 5e-2 1e-3 5e-2 5e-2 5e-2 5e-2 5e-2
@@
Hyper-parameter AdamW AdaMuon Muon+ (colrow) Mousse Muon-Nes MuonEq-Nes (RC) MuonEq-Nes (C) MuonEq-Nes (R)
Max Learning Rate 2e-3 5e-3 5e-3 2e-3 2e-3 2e-3 2e-3 2e-3
@@
We compare the three MuonEq-Nes variants RC, R, and C. RC applies two-sided row/column normalization throughout training, R applies row normalization throughout training, and C applies column normalization throughout training. We take R as the default variant. In all cases, the orthogonalization routine, optimizer states, and scalar hyperparameters are unchanged.
@@
This ablation is directly connected to Eq. (5). RC provides the strongest two-sided spectral correction and mainly targets the finite-step Newton–Schulz term. R and C test the two one-sided geometries under the same optimizer configuration. Since MuonEq/MuonEq-Nes targets hidden matrix weights, R is the more relevant default in our setting, while C is included as the column-sided companion form.
K.3 A Note on Row/Column Terminology in Prior Work
A careful comparison to prior row/column normalization papers requires fixing the matrix convention. [14] write linear weights as $\theta\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ and define column-wise normalization along the output dimension. Under the standard deep-learning storage layout $\mathbf{W}=\theta^{\top}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$, this corresponds to row normalization of the stored hidden-weight matrix. Their empirical observation that row-wise normalization can be unstable is also traced largely to the LM-head, where row-wise normalization produces extreme gradient values. This is not the same setting as MuonEq/MuonEq-Nes, which is aimed at hidden matrix weights.
@@
Appendix L Limitations
MuonEq is designed as a lightweight pre-orthogonalization correction for matrix-valued hidden weights. The row-normalized variant is our default in this setting, but it should not be viewed as a universally optimal choice for every parameter block or architecture. Our analysis isolates the core geometry of the pre-balanced orthogonalized update, while broader coverage of implementation variants and training regimes remains an open direction. Our convergence analysis for the RC variant is restricted to an auxiliary exact-polar regime: Proposition 3.4 assumes Nesterov momentum is disabled, no decoupled weight decay is used, and the stabilization parameter $\varepsilon$ is sufficiently large. Therefore, this RC guarantee does not formally cover the practical RC ablations, which use finite-step Newton–Schulz iterations and the small numerical stabilization $\varepsilon=10^{-8}$. Extending the RC theory to this practical regime is left for future work. This limitation is specific to the auxiliary RC analysis; the main guarantee for the default R variant is given by Theorem 3.5 and Corollary 3.6. Empirically, our large-scale evaluation focuses on LLaMA-style pretraining on C4, complemented by smaller-scale ablations.
Appendix M Broader Impacts
This work studies an application-agnostic optimizer that may improve training efficiency for matrix-valued neural network parameters without introducing new data, data collection, or deployment pipelines. Its direct societal risks are therefore mainly inherited from the downstream models and datasets on which it is used. As with other efficiency improvements, MuonEq could lower the cost of both beneficial and harmful model training, so appropriate downstream safeguards remain necessary. Any environmental benefit depends on whether efficiency gains are used to reduce compute for a fixed objective rather than to scale up training.